# T-cell receptor specificity landscape revealed through de novo peptide design

### v1

@@ IV Discussion
We found that HERMES-predicted peptide energies closely tracked both TCR-pMHC binding affinities and peptide-induced T-cell activities across diverse systems, including variants from a CMV-derived viral epitope [36], the NY-ESO cancer-testis antigen [43, 44], and a melanoma self-antigen [36]. Notably, HERMES also accurately predicted the activity of p53-derived neoepitope variants in the context of a bispecific antibody-mediated T-cell therapy [5]. This result underscores the algorithm’s broad applicability and generalizability: given a reliable structural template close to the system of interest, HERMES can predict peptide-induced T-cell responses, whether through direct TCR interactions or antibody-mediated mechanisms.
@@
Based on the HERMES predictions for TCR-pMHC reactivities, we presented a computational protocol for de novo design of immunogenic peptides that can elicit robust T-cell responses. Starting from the structural template of a TCR-MHC with a native (wild-type) peptide, we developed an MCMC search guided by HERMES peptide energies to explore the vast sequence space of roughly $10^{8}$ ($10^{10}$) possibilities for 9-residue (11-residue) peptides, to identify immunogenic candidates. Experimental validation across three systems—with native peptides from NY-ESO and MAGE cancer epitopes, plus a viral epitope from EBV—revealed that our designs reliably activated T-cells, even with up to five substitutions from the native peptide used in the structural template. Notably, several MAGE-derived variants outperformed or matched the native peptide in T-cell activation, demonstrating the method’s potential to enhance peptide potency. These results show the potential of HERMES for designing peptide-based cancer vaccines with neoantigens, providing an avenue for personalized immunotherapy strategies.
@@
Accurately modeling T-cell specificity remains challenging, in part due to the limited functional data on TCR-epitope pairs, and the highly cross-reactive nature of TCR-epitope interactions. A design protocol like the one presented here, coupled with experimental validation, offers a promising path forward for future studies. By iteratively generating candidate peptide libraries, testing them experimentally, and refining the model with each iteration, it should be possible to more efficiently explore the high dimensional antigenic sequence space. This active learning strategy has a potential for revealing the “shape space” of TCR-epitope interactions, with broad applicabilities for early disease diagnosis, and for design of targeted immunotherapies. Moreover, this approach can help mitigate immunotherapy-related toxicity by identifying and excluding potential self-reactivities–beyond single point mutations in the target epitope–such as the off-target Titin reactivity observed in an engineered T-cell therapy against MAGE-A3 that led to a severe cardiac toxicity [57].
@@
Although our study primarily examines MHC class I-restricted peptides, the broad applicability of structure-based modeling suggests that similar methods could be adapted for the more complex TCR-peptide-MHCII interactions, thereby offering new insights into CD4+ T-cell responses. Moreover, incorporating additional immunological factors—such as TCR clustering, co-receptor interactions, cross-reactivity with self-antigens, and functional immune profiling—can further enhance the accuracy of T-cell response predictions and improve the resulting de novo peptide design pipeline.
V Data and code availability

### v2

@@ Discussion
We found that HERMES-predicted peptide energies closely tracked both TCR-pMHC binding affinities and peptide-induced T-cell activities across diverse systems, including variants from a CMV-derived viral epitope [52], the NY-ESO cancer-testis antigen [48, 49], and a melanoma self-antigen [52]. Notably, HERMES also accurately predicted the activity of p53-derived neoepitope variants in the context of a bispecific antibody-mediated T-cell therapy [5]. This result suggests the algorithm’s potential applicability: given a reliable structural template close to the system of interest, HERMES can predict peptide-induced T-cell responses, whether through direct TCR interactions or antibody-mediated mechanisms. However, drawing more definitive conclusions about the generalizability of HERMES would require examining a larger datasets of TCRs interacting with pMHC variants.
@@
Based on the HERMES predictions for TCR-pMHC reactivities, we presented a computational protocol for de novo design of immunogenic peptides that can elicit robust T-cell responses. Starting from the structural template of a TCR-MHC with a native (wild-type) peptide, we developed an MCMC search guided by HERMES peptide energies to explore the vast sequence space of roughly $10^{8}$ ($10^{10}$) possibilities for 9-residue (11-residue) peptides, to identify immunogenic candidates. Experimental validation across three systems—with native peptides from NY-ESO and MAGE cancer epitopes, plus a viral epitope from EBV—revealed that our designs reliably activated T-cells, even with up to five substitutions from the native peptide used in the structural template. Notably, several MAGE-derived variants outperformed or matched the native peptide in T-cell activation, demonstrating the method’s potential to enhance peptide potency.
This capability can be relevant to personalized peptide-based cancer vaccines, which aim to induce durable T-cell memory against a patient’s tumor antigens to reduce the risk of rebound [6, 69, 8]. Despite their promise, the success of cancer vaccines has been limited by low peptide immunogenicity and poor uptake by antigen-presenting cells [70]. Recent approaches have employed deep mutational scanning experiments to identify more immunogenic cancer neoantigens against specific TCR candidates [52]. However, these experiments are often limited to scanning single amino acid mutations from the wild-type peptide due to the combinatorial explosion when exploring multiple mutations. Moreover, the precise TCRs responding to peptide vaccines in a patient are often unknown, and many immune responses are induced de novo [69]. With the caveat of access to suitable structural templates, HERMES could inform T-cell vaccine design by computationally generating diverse peptide libraries predicted to engage a broad set of candidate TCRs (identified independently), nominating vaccine epitopes for subsequent safety and efficacy testing.
@@
Accurately modeling T-cell specificity remains challenging, in part due to the limited functional data on TCR-epitope pairs, and the highly cross-reactive nature of TCR-epitope interactions. A design protocol like the one presented here, coupled with experimental validation, offers a promising path forward for future studies. By iteratively generating candidate peptide libraries, testing them experimentally, and refining the model with each iteration, it should be possible to more efficiently explore the high dimensional antigenic sequence space. This active learning strategy has a potential for revealing the “shape space” of TCR-epitope interactions, with broad applicabilities for early disease diagnosis, and for design of targeted immunotherapies. Moreover, this approach can help mitigate immunotherapy-related toxicity by identifying and excluding potential self-reactivities–beyond single-point mutations in the target epitope–such as the off-target Titin reactivity observed in an engineered T-cell therapy against MAGE-A3 that led to a severe cardiac toxicity [59].
@@
TCRs can cross-react with diverse peptides in multiple orientations and conformation [65]. To reflect this flexibility, we sampled peptide conformational space (i) implicitly, by applying noise to the coordinates during training, and (ii) explicitly, by locally relaxing the atomic environment after each amino acid substitution. Even so, AlphaFold3 predicts that our experimentally validated immunogenic designed peptides retain similar structural poses as their wild-type templates. This structural convergence may partly stem from AlphaFold3’s bias toward structures represented in its training set, but it also exposes limitations in our current sampling strategy. Future work should therefore aim to co-design peptide sequence, peptide backbone geometry, together with the CDR loop conformations of the interacting TCR, enabling a more comprehensive exploration of the peptide-TCR conformational landscape.
Although our study primarily examines MHC class I-restricted peptides, the broad applicability of structure-based modeling suggests that similar methods could be adapted for the more complex TCR-peptide-MHCII interactions, thereby offering new insights into CD4+ T-cell responses. Moreover, incorporating additional immunological factors—such as TCR clustering, co-receptor interactions, cross-reactivity with self-antigens, and functional immune profiling—can further enhance the accuracy of T-cell response predictions and improve the resulting de novo peptide desig pipeline.
II Data and code availability